OpenStudy (anonymous):

Can anyone help stuck on differentiation question y= 5 sin 2x

OpenStudy (anonymous):

I guess that you have to differentiate function \[f(x)=5sin2x\] So, you have to use two differentiation rules: constant rule and chain rule. The first is because you have 5 as a constant. This constant is factored out and it doesn't play role in the main part of work. The main part of work uses chain rule on \[sin2x\] To use it we have to substitute 2x by \[u\] and this gives us \[sin u\] Now, use the chain rule \[{df \over dx} = {df \over du} * {du \over dx}\]

OpenStudy (anonymous):

Thanks will try it

OpenStudy (anonymous):

You're welcome. Remember to multiply by factored constant in the end.

OpenStudy (anonymous):

3x+(-7)=-11+(-11)

OpenStudy (anonymous):

3x+(-7)=-11+(-11)

OpenStudy (anonymous):

answer this plsss

OpenStudy (anonymous):

Well, \[3x - 7 = -22\] then add seven to both sides \[3x = -15\] Finally divide both sides by three \[x = -5\]